in fall 2011 the euro dollar exchange rate was $1.35 and by spring 2015 it strenghthed to $1.10 . Asssume that the euorpean luxury marketeres cut the price of an $8000 linen suit by 10 percent when launching its spring 2015 collection . how would revenues have been affected when dollar prices were converted to euros?
Answer #1

Initial price in euro = 8000 x 1.35 = 10800

Spring price in euro = 8000 x 90% x 1.1 = 7920

Assuming quantity remains unchanged,

Change in revenue = Change in euro price

Change in euro price (in euro) = 7920 - 10800 = - 2880 (decrease)

Change in euro price (in %) = (7920 / 10800) - 1 = 0.7333 - 1 = - 0.2667 = - 26.67%